The bulk modulus of a block of copper is 1.3 GPa. When a pressure of 1.8 MPa is applied to the 1.0 m^3 block, its volume decreases by how much?
1
Expert's answer
2020-02-12T10:41:45-0500

"K=-V_0\frac{\Delta p}{\Delta V}"


Given.


"K=1.3 GPa"


"p_0=101325 Pa"


"p=1.8 MPa"


"V_0=1 m^3"


"\Delta V=-V_0\frac{p-p_0}{K}=-1\cdot \frac{1800000-101325}{1.3\cdot10^9}\approx-0.00131 m^3=-1310 cm^3"